# Understanding and Mitigating Tokenization Bias in Language Models

### v1

@@ E.1 Refactoring
where the last inequality is due to Corollary 3.2. Finally, we will use the Branch and Pass Algorithm to compute each term in (14) individually. Note that the algorithm does not require $t_{k}\in\mathcal{V}^{*}$. Here, we explicitly highlight the importance of having $t_{k}\in\mathcal{V}^{*}$, as it bridges between the character and token domain through Equation (14).
E.2 The Branch and Pass Algorithm
Overview. The Branch and Pass (BPTree) algorithm has two inputs: the query string value $x^{N}_{n_{k}+1}$ and the conditioning tokens $t^{k}_{1}$, which outputs the value $P(x^{N}_{n_{k}+1}|t^{k}_{1})$. Note that we do not require $t_{k}\in\mathcal{V}^{*}$ in the BPTree algorithm. First, using marginalization, we have the following:
@@ E.2 The Branch and Pass Algorithm
where $\mathcal{T}_{\mathrm{b_{val}}}{=}\{t|t{\in}\mathcal{V}{\wedge}x^{N}_{n_{k}+1}{\in}\mathrm{prefix}(\mathrm{decode}(t))\}$ and $\mathcal{T}_{\mathrm{p_{val}}}{=}\{t|t\in\mathcal{V}\wedge x^{N}_{n_{k}+1}{\notin}\mathrm{prefix}(\mathrm{decode}(t))\}$, where $\mathcal{T}_{\mathrm{b_{val}}}\cap\mathcal{T}_{\mathrm{p_{val}}}=\emptyset$, are the corresponding set we iterates over each sum in the branch step, i.e. $\mathrm{b_{val}}$, and pass step, i.e. $\mathrm{p_{val}}$, respectively. Intuitvely, $\mathcal{T}_{\mathrm{b_{val}}}$ is the set of tokens that have a prefix $x^{N}_{n_{k}+1}$ and $\mathcal{T}_{\mathrm{p_{val}}}$ is the ones that do not. We now show how to compute each term separately using a tokenized LM that outputs $P(t_{i+1}|t^{i}_{1})$.
Branch Step. Note that $\mathrm{b_{val}}$ can be described as the probability that, given the list of previous tokens $t^{k}_{1}$, the next token of the string $s$ has $x^{N}_{n_{k}+1}$ as a prefix. As such, to obtain $\mathrm{b_{val}}$, we first compute $P(t_{k+1}=t|t^{k}_{1})$ for all $t\in\mathcal{V}$ using one model run, then sum the probabilities corresponds to all tokens whose prefix is $x^{N}_{n_{k}+1}$. Specifically:
@@
where the second equation is due to $x^{N}_{n_{k}+1}{\in}\mathrm{prefix}(t)$ so $P(x^{N}_{n_{k}+1}|t^{k}_{1},t_{k+1}=t){=}1.0$. This concludes the proof. ∎
Pass Step. In contrast to $\mathrm{b_{val}}$, the $\mathrm{p_{val}}$ is the probability that, given the list of previous (valid) tokens $t^{k}_{1}$, the string $s$ has the subsequent string as $x^{N}_{n_{k}+1}$ that is not a prefix of the next token. We will show that under the maximum prefix encoding, we compute the value $\mathrm{p_{val}}$ as follow:
@@
    Obtain the next token $t$ within $x^{N}_{n_{k}+1}$ and compute $P(t_{k+1}=t|t^{k}_{1})$. If $x^{N}_{n_{k}+1}=\mathrm{decode}(t)$, then returns $0.0$ since this is not allowed according to the condition required in $\mathcal{T}_{\mathrm{p_{val}}}$.
@@
We first rewrite $\mathcal{T}_{\mathrm{p_{val}}}$ as the union of the following disjoint sets, i.e $\mathcal{T}_{\mathrm{p_{val}}}{=}\mathcal{T}_{\mathrm{p_{val,1}}}{\cup}\mathcal{T}_{\mathrm{p_{val,2}}}$ and $\mathcal{T}_{\mathrm{p_{val,1}}}{\cap}\mathcal{T}_{\mathrm{p_{val,2}}}{=}\emptyset$, where:
  • $\mathcal{T}_{\mathrm{p_{val,1}}}=\{t|t\in\mathcal{V}\wedge x^{N}_{n_{k}+1}\notin\mathrm{prefix}(\mathrm{decode}(t))\wedge\mathrm{decode}(t)\notin\mathrm{prefix}(x^{N}_{n_{k}+1})\}$: the subset of $\mathcal{T}_{\mathrm{p_{val}}}$ where the string represented by the token $t$ is not a prefix of $x^{N}_{n_{k}+1}$.
  • $\mathcal{T}_{\mathrm{p_{val,2}}}=\{t|t\in\mathcal{V}\wedge x^{N}_{n_{k}+1}\notin\mathrm{prefix}(\mathrm{decode}(t))\wedge\mathrm{decode}(t)\in\mathrm{prefix}(x^{N}_{n_{k}+1})\}$: the subset of $\mathcal{T}_{\mathrm{p_{val}}}$ where string represented by the token $t$ is actually a prefix of $x^{N}_{n_{k}+1}$.
As such, we have:
$\mathrm{p_{val}}=\sum_{t\in\mathcal{T}_{\mathrm{p_{val}}}}P(x^{N}_{n_{k}+1},t_{k+1}=t|t^{k}_{1})=\underbrace{\sum_{t\in\mathcal{T}_{\mathrm{p_{val,1}}}}P(x^{N}_{n_{k}+1},t_{k+1}=t|t^{k}_{1})}_{0.0}+\sum_{t\in\mathcal{T}_{\mathrm{p_{val,2}}}}P(x^{N}_{n_{k}+1},t_{k+1}=t|t^{k}_{1}),$ (21)
where we note that for $t\in\mathcal{T}_{\mathrm{p_{val,1}}}$, we have $P(x^{N}_{n_{k}+1}|t^{k}_{1},t_{k+1}=t)=0.0$ since the conditional $t_{k+1}=t$ implies that the string represented by $t$ must be a prefix of $x^{N}_{n_{k}+1}$, which leads to the first sum being $0.0$.
For the second term, i.e. $t\in\mathcal{T}_{\mathrm{p_{val,2}}}$, if $t\neq\mathrm{encode}(x^{N}_{n_{k}+1})_{1}$, by following Proposition 2.3, we have $P(x^{N}_{n_{k}+1},t_{k+1}=t|t^{k}_{1})=0.0$ since any tokenization that is not a valid maximum prefix encoding must happen with probability $0.0$. This stands corrected as long as the output probability of our LM follows Proposition 2.3 (see Section G.2).
As a result, we are left with the token $t=\mathrm{encode}(x^{N}_{n_{k}+1})_{1}$, which is the Equation (20) after applying chain rule of probability. ∎
@@
Appendix F Converting Token-Free Language Model to Tokenized Language Model.
@@ F.1 Refactoring
This concludes the proof.
@@
Appendix G On Predictive Distribution of LMs
In practice, our LM does not follow Proposition 2.3 due to softmax activations. As such, in our BPTree algorithm, when $t\in\mathcal{T}_{\mathrm{p_{val,2}}}$ and $t\neq\mathrm{encode}(x^{N}_{n_{k}+1})_{1}$, then $P_{\theta}(x^{N}_{n_{k}+1},t_{k+1}=t|t^{k}_{1})$ may not be $0.0$ (where $\theta$ is the model weights). Eventually, this can potentially increase the complexity of our BPTree algorithm during the Passing step.
In the first part of this section, we show that given any tokenized LM, we can force its output probabilites to obey Proposition 2.3, without any loss in terms of perplexity score on the token domain. In the second part, for completeness, we will show that having a tokenized LM satifying Propostion 2.3 will guarantee the correctness of the Passing step in our BPTree algorithm.
@@ G.1 Truncate-Renormalization Process
We justify the assumption that our tokenized language model $P_{\theta}(t_{i+1}|t^{i}_{1})$ follows Proposition 2.3. The idea is that we can turn a language model that does not follow Proposition 2.3 to the one that does while guaranteeing that the new model will always results in a higher token-level perplexity score.
@@
which implies that $q^{*}$ is closer to $p$ than $q$. We refer the process of producing $q^{*}$ as truncate-renormalization (TR).
@@
Let $Z=(\sum_{l\notin\Phi}q_{l})$ is the normalizing factor in $q^{*}$. Note that $Z\leq 1$ and as such $\log(Z)\leq 0$. The proof is as follows:
@@
Applying this to our scenario, where we are given any autoregressive language models $\hat{P}_{\theta}(t_{i+1}|t^{i}_{1})$ that does not follow Proposition 2.3 (due to the softmax activations), we can perform the TR process (since we know which encoding is invalid) to obtain a new distribution ${P}_{\theta}(t_{i+1}|t^{i}_{1})$, which is guaranteed to be closer to the ground-truth model ${P_{\mathrm{gt}}}(t_{i+1}|t^{i}_{1})$. As the results, we are guaranteed that the token-level perplexity score of ${P}_{\theta}(t_{i+1}|t^{i}_{1})$ is always lower than or equal to $\hat{P}_{\theta}(t_{i+1}|t^{i}_{1})$.
G.2 On Passing Step in Branch and Pass Algorithm.
Once our tokenized LM follows Proposition 2.3, intuitively it can be regarded as an autogressive source (in the token domain) so it does not alternate the correctness of the Passing step. This section provides a proof for completeness.
Note that we only need to consider the second term of $\mathrm{p_{val}}$, i.e.:
$\mathrm{p_{val}}=\underbrace{\sum_{t\in\mathcal{T}_{\mathrm{p_{val,1}}}}P(x^{N}_{n_{k}+1},t_{k+1}=t|t^{k}_{1})}_{0.0}+\sum_{t\in\mathcal{T}_{\mathrm{p_{val,2}}}}P(x^{N}_{n_{k}+1},t_{k+1}=t|t^{k}_{1}),$ (50)
as the other term is always $0.0$ regardless. Specifically, we will prove that, during the Pass step, when $t\in\mathcal{T}_{\mathrm{p_{val,2}}}$ and $t\neq\mathrm{encode}(x^{N}_{n_{k}+1})_{1}$, then $P(x^{N}_{n_{k}+1},t_{k+1}=t|t^{k}_{1})=0.0$ given that our LM follows Proposition 2.3. If $t^{k+1}_{1}$ is an invalid encoding, we immediately have $P(x^{N}_{n_{k}+1},t_{k+1}=t|t^{k}_{1})=0.0$, so we only consider when $t^{k+1}_{1}$ is a valid encoding. We first prove the following Proposition.
Proposition G.2.
If $t^{k}_{1}$ is invalid, then $t^{k+1}_{1}$ is also invalid.
@@
Proof by contradiction, suppose $t^{k+1}_{1}$ is valid then $\mathcal{S}(t^{k+1}_{1})\neq\emptyset$. However, any string $s\in\mathcal{S}(t^{k+1}_{1})$ must also be in $\mathcal{S}(t^{k}_{1})$. But $\mathcal{S}(t^{k+1}_{1})=\emptyset$ (Corollary B.3). As such, $t^{k+1}_{1}$ is invalid. ∎
Since $x^{N}_{n_{k}+1}$ has totally $m=N-(n_{k}+1)$ characters, it can have maximum $m$ tokens. Let $\mathcal{C}=\{\tau\in\mathcal{V}^{m-1}|x^{N}_{n_{k}+2}\in\mathrm{prefix}(\mathrm{decode}(\tau))\}$ be the set of all possible sequence of $m-1$ tokens (both valid and invalid). We then have:
$P(x^{N}_{n_{k}+1},t_{k+1}=t|t^{k}_{1})$ $=\sum_{\tau\in\mathcal{C}}P(x^{N}_{n_{k}+1},t_{k+1}=t,t^{k+m}_{k+2}=\tau|t^{k}_{1})$ (51)
$=\sum_{\tau\in\mathcal{C}}P(t_{k+1}=t,t^{k+m}_{k+2}=\tau|t^{k}_{1})P(x^{N}_{n_{k}+1}|t^{k}_{1},t_{k+1}=t,t^{m}_{k+2}=\tau)$ (52)
Then, it is sufficient to show that $P(t_{k+1}=t,t^{k+m}_{k+2}=\tau|t^{k}_{1})=0.0$ for any $\tau\in\mathcal{C}$ as long as $t$ does not follow the maximum prefix encoding, i.e. $t\neq\mathrm{encode}(x^{N}_{n_{k}+1})_{1}$. We have the followings:
  • Since $t_{k+1}=t$ does not follow the maximum prefix encoding, then $\mathrm{encode}(\mathrm{decode}(t^{k+m}_{1})){\neq}t^{k+m}_{1}$ since we can always change the tokenization at $t_{k+1}$.
  • Let $t_{k+j}$ be the first token where $t^{k+j}_{1}$ is invalid. Then follow Proposition G.2, since $t^{k+j}_{1}$ is invalid, then any extra tokens after $t^{k+j}_{1}$ must also be invalid.
As such, we have: $P(t_{k+j}|t^{k+j-1}_{1})=0.0$, which leads to $P(t_{k+1}=t,t^{k+m}_{k+2}=\tau|t^{k}_{1})=0.0$, which proves the results.

### v2

@@ E.1 Refactoring
where the last inequality is due to Corollary 3.2. Finally, we will use the Maximum Prefix Correction (MPE) Algorithm to compute each term in (14) individually. Note that the algorithm does not require $t_{k}\in\mathcal{V}^{*}$. Here, we explicitly highlight the importance of having $t_{k}\in\mathcal{V}^{*}$, as it bridges between the character and token domain through Equation (14).
E.2 Maximum Prefix Correction Algorithm
Overview. The MPC algorithm computes $P(x^{N}_{n_{k}+1}|t^{k}_{1})$. Note that we do not require $t_{k}\in\mathcal{V}^{*}$ in the MPC algorithm. Using marginalization, we have the following:
@@ E.2 Maximum Prefix Correction Algorithm
where:
  • $\mathcal{T}_{\mathrm{b_{val}}}=\{t\in\mathcal{V}|x^{N}_{n_{k}+1}\in\mathrm{prefix}(\mathrm{decode}(t))\}$ is the set of tokens that have a prefix $x^{N}_{n_{k}+1}$.
  • $\mathcal{T}_{\mathrm{p_{val}}}=\{t\in\mathcal{V}|x^{N}_{n_{k}+1}\notin\mathrm{prefix}(\mathrm{decode}(t))\}$ is the ones that do not.
and $\mathcal{T}_{\mathrm{b_{val}}}\cap\mathcal{T}_{\mathrm{p_{val}}}=\emptyset$.
Branch Step. Here, $\mathrm{b_{val}}$ is the probability that, given the list of previous tokens $t^{k}_{1}$, the next token of the string $s$ has $x^{N}_{n_{k}+1}$ as a prefix. To compute this term, we obtain $P(t_{k+1}=t|t^{k}_{1})$ for all $t\in\mathcal{V}$ using one model run, then sum the probabilities corresponds to all tokens whose prefix is $x^{N}_{n_{k}+1}$.
@@
where $P(x^{N}_{n_{k}+1}|t^{k}_{1},t_{k+1}{=}t)=1.0$ is due to $x^{N}_{n_{k}+1}\in\mathrm{prefix}(t)$. This concludes the proof. ∎
Pass Step. Here, $\mathrm{p_{val}}$ is the probability that, given the list of previous tokens $t^{k}_{1}$, the subsequent string $x^{N}_{n_{k}+1}$ is not a prefix of the next token. Under the MPE, we compute the value $\mathrm{p_{val}}$ as follow:
@@
    Extract the next token $t$ within $x^{N}_{n_{k}+1}$ and compute $P(t_{k+1}=t|t^{k}_{1})$. If $x^{N}_{n_{k}+1}=\mathrm{decode}(t)$, then returns $0.0$ since this is not allowed according to the condition required in $\mathcal{T}_{\mathrm{p_{val}}}$.
@@
Following Proposition 2.3 for invalid encodings, we only need to consider $t$ such that $t^{k+1}_{1}$ is valid. Under Proposition B.1 for MPE on $x^{N}_{1}$, only first token of $\mathrm{encode}(x^{N}_{n_{k}+1})$ is allowed (also see Example 2 in Figure 4(Right)). Finally, applying the chain rule of probability, we obtain Equation 20. For the case of non-optimal LM, see Section G.2 for non-optimal LM. This completes the proof. ∎
@@
Appendix F Converting Token-Free Language Model to Tokenized Language Model for MPE.
@@ F.1 Refactoring
This concludes the proof. ∎
@@
Appendix G On Predictive Distribution of Language Models
In practice, LMs often do not follow Proposition 2.3 due to softmax activations. As such, in our MPC algorithm, when $t\in\mathcal{T}_{\mathrm{p_{val}}}$ and $t\neq\mathrm{encode}(x^{N}_{n_{k}+1})_{1}$, then $P_{\theta}(x^{N}_{n_{k}+1},t_{k+1}=t|t^{k}_{1})$ may not be $0.0$ (where $\theta$ is the model weights). Eventually, this can potentially increase the complexity of our MPC algorithm during the Passing step.
In this section, we show that given any tokenized LM, we can force its output probabilities to obey Proposition 2.3, without any loss in terms of perplexity score on the token domain. This means that a tokenized LM satisfying Proposition 2.3 will guarantee the correctness of the Passing step in our MPC algorithm.
@@ G.1 Truncate-Renormalization Process
We justify the assumption that our tokenized language model $P_{\theta}(t_{i+1}|t^{i}_{1})$ follows Proposition 2.3. The idea is that we can turn a language model that does not follow Proposition 2.3 to the one that does while guaranteeing that the new model will always result in a lower token-level perplexity score.
@@
which implies that $q^{*}$ is closer to $p$ than $q$. We refer to the process of producing $q^{*}$ as truncate-renormalization (TR).
@@
Let $Z=(\sum_{l\notin\Phi}q_{l})$ is the normalizing factor in $q^{*}$. Note that $Z\leq 1$ and as such $\log(Z)\leq 0$. Then:
@@
Applying to our scenario, for any autoregressive language models $\hat{P}_{\theta}(t_{i+1}|t^{i}_{1})$ that does not follow Proposition 2.3 (due to the softmax activations), we can perform the TR process (since we know which encoding is invalid) to obtain a new LM ${P}_{\theta}(t_{i+1}|t^{i}_{1})$, which is guaranteed to better approximate the ground-truth model ${P_{\mathrm{gt}}}(t_{i+1}|t^{i}_{1})$. Thus, we are guaranteed that the token-level perplexity score of ${P}_{\theta}(t_{i+1}|t^{i}_{1})$ is always lower than or equal to $\hat{P}_{\theta}(t_{i+1}|t^{i}_{1})$.
G.2 On Passing Step in Maximum Prefix Correction Algorithm.
Once our tokenized LM follows Proposition 2.3, it does not alternate the correctness of the Passing step. In other words, under Proposition 2.3, the LM will always output zero probability for invalid encodings $t^{k}_{1}$. As a result, the Passing step in the MPC algorithm remains the same in this case.
Appendix H Algorithms for Byte Pair Encoding
H.1 Overview
We begin by introducing the Byte-Pair Correction (BPC) Algorithm for bias correction in Byte-Pair Encoding, which is more general than the MPC algorithm and also works for case of MPE. We then follow with a detail analysis to show the correctness of the algorithm.
Here, we introduce the definitions of invalid encodings (for BPE) and cover encodings.
Definition H.1.
(Invalid Encodings) The list of tokens (an encoding) $t^{k}_{1}$ is invalid if $\mathrm{encode}(\mathrm{decode}(t^{k}_{1})){\neq}t^{k}_{1}$. Otherwise, it is a valid encoding. We denote a valid $t^{k}_{1}$ as $\mathrm{valid}(t^{k}_{1})$.
Definition H.2.
(Cover Encodings) Given a string $x^{n}_{1}$, an encoding $t^{k}_{1}$ is said to be covering $x^{n}_{1}$ when all the following conditions satisfied:
  1. 1.
    $t^{k}_{1}$ is valid.
  2. 2.
    $x^{n}_{1}\in\mathrm{prefix}(\mathrm{decode}(t^{k}_{1}))$.
  3. 3.
    $x^{n}_{i}\in t_{k}$ for some $1\leq i\leq n$, i.e. the last token $t_{i}$ covers a part of the string $x^{n}_{1}$.
We denote $\mathrm{cover}(x^{n}_{1})$ to be the set of all cover encodings of $x^{n}_{1}$ and $\vec{t}\in\mathrm{cover}(x^{n}_{1})$ is an encoding in $\mathrm{cover}(x^{n}_{1})$.
Having established these two definitions, we will later show that for BPE (and MPE), the probability $P(x^{n}_{1})$ can be represented using a tokenized LM $P(t_{i+1}|t^{i}_{1})$ as follows:
$P(x^{n}_{1})=\sum_{\vec{t}\in\mathrm{cover}(x^{n}_{1})}P(\vec{t}),$ (49)
and the main goal of the BPC algorithm is to search through all cover encodings of $x^{n}_{1}$. We can then apply this algorithm and compute any conditional probability $P(x^{N}_{n+1}|x^{n}_{1})$ through factorization. Figure 5 (Left) illustrates this with examples cover encodings and invalid/valid encodings.
H.2 Byte-Pair Correction Algorithm
Figure 5: (Left) Representation of $P(x^{n}_{1})$ using tokenized LM. We also show an example of cover encodings and valid/invalid encoding. (Right) Illustration of the Byte-Pair Correction Algorithm for BPE encoding. Green tick and red cross denotes valid and invalid encodings respectively, which can be checked using Definition H.1. The termination step does not appear in the original algorithm (for simplicity) but can be easily implemented.
[…]
Algorithm 2 Byte-Pair Correction Algorithm. This algorithm computes $P(x^{n}_{1})$ by gradually reducing the search space.
1:procedure compute($x^{N}_{1}$)
2:    //Initialize $P(x^{n}_{1})$:
3:    $P_{out}=0.0$
4:    //Probability Aggregation
5:    for $i=n-1...0$ do
6:        //The last token that partially covers $x^{n}_{1}$ begins with $x^{n}_{i+1}$
7:        =$\mathcal{B}=\{t\in\mathcal{V}|x^{n}_{i{+}1}{\in}\mathrm{prefix}(\mathrm{decode}({t}))\}$
8:        //Once the last token position is known, the remaining previous tokens can be determined
9:        $t^{k-1}_{1}=\mathrm{encode}(x^{i}_{1})$
10:        //Similar to the branching step in MPC
11:        $\mathrm{b_{val}}={\sum\limits_{t{\in}\mathcal{B}}}P(t_{k}=t\big{|}t^{k-1}_{1})$
12:        $P_{out}=P_{out}+P(t^{k-1}_{1})\times\mathrm{b_{val}}$
13:    end for
14:return $P_{out}$
15:end procedure
The Byte-Pair Correction (BPC) algorithm, shown in Algorithm 2 and visualized in Figure 5 (right), which is an efficient algorithm that can search all valid encodings covering $x^{n}_{1}$. The idea is that, for each cover encoding $\vec{t}$, once the starting position of the last token is determined (say $x_{i+1}$), we are guaranteed the prior tokens is unique and must be $\mathrm{encode}(x^{i}_{1})$. Then one will accept the extracted $\vec{t}$ if it is valid, otherwise reject it. Corollary H.3 provides justification for this procedure. Here, we assume $P(\vec{t})=0.0$ for invalid $\vec{t}$, see Proposition H.4 and justifications as well as implementation in Appendix G.
Remark. The BPC algorithm can also be applied for the case of MPE. In fact, it is more general than the original MPC algorithm as it only relies on the property of invalid encodings.
H.3 Analysis
Notations. We extend the notation of the vocabulary $\mathcal{V}$ for the case of BPE. Here, $\mathcal{V}$ is an ordered list that determines the merging order in the BPE algorithm. Each $v\in\mathcal{V}$ is a triplet of $(t_{\mathrm{left}},t_{\mathrm{right}},t_{\mathrm{new}})$ which corresponds to the merging tokens (left and right) and the new token. For simplicity, when we write $t\in\mathcal{V}$, it corresponds to the merged token, i.e. $t_{\mathrm{new}}\in v$. Finally, the first $|\mathcal{A}|$ entries in $\mathcal{V}$ correspond to the alphabet $\mathcal{A}$, where no merging will happen.
Byte-Pair Encoding. We revise the encoding rule for BPE, shown in Algorithm 3. In practice, pre-tokenization is often used, where tokens are separated by whitespace or special characters. In this case, we can adjust our vocabulary $\mathcal{V}$ by removing tokens with special characters in the middle of the string.
Algorithm 3 Byte Pair Encoding Algorithm .
1:procedure Encode_BPE($x^{N}_{1}$, $\mathcal{V}$)
2:    //Set initial encodings:
3:    $\mathrm{c\_tokens}=x^{N}_{1}$
4:    //Iterate over merging order in $\mathcal{V}$, the first $|\mathcal{A}|$ entries correspond the the alphabet (no merge happens):
5:    for $i=|\mathcal{A}|+1,...|\mathcal{V}|$ do
6:        $\mathrm{c\_tokens}\xleftarrow{}\mathrm{find\_merge}(\mathrm{c\_tokens},\mathcal{V}[i])$
7:    end for
8:    return $\mathrm{c\_tokens}$
9:end procedure
10:
11:procedure $\mathrm{find\_merge}$($\mathrm{c\_tokens},v$)
12:    // Left and right tokens for merging
13:    $t_{\mathrm{left}},t_{\mathrm{right}},t_{\mathrm{new}}=v[1],v[2],v[3]$
14:    $\mathrm{old\_tokens}=\mathrm{c\_tokens}$
15:    $\mathrm{new\_tokens}=[]$
16:    // Find and merge tokens from left to right
17:    $j=1$
18:    while $j<|\mathrm{old\_tokens}|$ do
19:        if $\mathrm{old\_tokens}[i,i+1]=t_{\mathrm{left}},t_{\mathrm{right}}$ then
20:           $\mathrm{new\_tokens.append}(t_{\mathrm{new}})$
21:           $j=j+2$
22:        else
23:           $\mathrm{new\_tokens.append}(\mathrm{old\_tokens}[i])$
24:           $j=j+1$
25:        end if
26:    end while
27:    return $\mathrm{new\_tokens}$
28:end procedure
Overview. We begin our analysis with theoretical results on invalid encodings (Corollary H.3 and Proposition H.4), which characterizes the zero probability events for an optimal tokenized LM. This will allow us to prove the representation of $P(x^{n}_{1})$, i.e. Proposition H.5, previously shown in Equation (49). Finally, we conclude this section with the proof of correctness of the BPC algorithm, using Proposition H.5 and H.6.
We begin with theoretical results on invalid encodings in the case of BPE.
Corollary H.3.
$\mathcal{S}(t^{k}_{1})=\emptyset$ if and only if $t^{k}_{1}$ is invalid.
@@ H.3 Analysis
We prove each direction as follows.
  • If $t^{k}_{1}$ is invalid then $\mathcal{S}(t^{k}_{1})=\emptyset$: Let $x^{n}_{1}=\mathrm{decode}(t^{k}_{1})$, we consider two string $s_{1}$ and $s_{2}$ that both have prefix $x^{n}_{1}$. Furthermore, we assume the first $i$ tokens of $s_{1}$ covers exactly $x^{n}_{1}$, i.e. $x^{n}_{1}=\mathrm{decode}(t^{i}_{1})$ and similarly, the first $j$ tokens of $s_{2}$ covers exactly $x^{n}_{1}$, i.e. $x^{n}_{1}=\mathrm{decode}(t^{j}_{1})$. Then:
    1. 1.
      Proving invalid $t^{k}_{1}$ leads to $\mathcal{S}(t^{k}_{1})=\emptyset$ is equivalently to proving $t^{i}_{1}=t^{j}_{1}$ for any $s_{1},s_{2}$.
    2. 2.
      Re-running the BPE algorithm for $s_{1}$ and $s_{2}$ in parallel, we know that there will be no merge between any suffix of $x^{n}_{1}$ and the rest of strings, i.e. $s_{1}\backslash x^{n}_{1}$ and $s_{2}\backslash x^{n}_{1}$ due to the condition above (See Algorithm 3, line 6).
    3. 3.
      Furthermore, for $s_{1}$, any time a merge happens within $x^{n}_{1}$ then the same merge must also happen within $x^{n}_{1}$ for $s_{2}$ and vice versa.
    As the result, we have $t^{i}_{1}=t^{j}_{1}$ and they must be equal to $\mathrm{encode}(x^{n}_{1})$.
Proposition H.4.
(Token-Induced Zero Probability-BPE) Let $t^{i}_{1}$ be a sequence of input tokens. For any invalid encoding $t^{i}_{1}$, we have $P_{\mathrm{gt}}(t^{i}_{1}){=}0.0$ and the conditional probability $P_{\mathrm{gt}}(t_{i+1}|t^{i}_{1})$ is undefined. In the case $t^{i}_{1}$ is valid, then $P_{\mathrm{gt}}(t_{i+1}|t^{i}_{1}){=}0.0$ if $t^{i+1}_{1}$ is invalid.
Proof.
The proof is the same as the MPE version (Proposition 2.3). ∎
Correctness of BPC Algorithm. We show in Proposition H.5 that computing the string probability $P(x^{n}_{1})$ is equivalent to marginalizing the probability of all covering tokens of $x^{n}_{1}$. As such, the main task of computing $P(x^{n}_{1})$ is to iterate all the valid encodings that cover $x^{n}_{1}$.
Proposition H.5.
(Prefix Probability Representation) Given a prefix $x^{n}_{1}$, we have the followings:
  1. 1.
    For any distinct $\vec{t}_{i},\vec{t}_{j}\in\mathrm{cover}(x^{n}_{1})$, then $\mathcal{S}(\vec{t}_{i})\cap\mathcal{S}(\vec{t}_{j}){=}\emptyset$.
  2. 2.
    $\mathcal{S}(x^{n}_{1})=\smashoperator[]{\bigcup_{\vec{t}\in\mathrm{cover}(x^{n}_{1})}^{}}\mathcal{S}(\vec{t})$.
As a result, $P(x^{n}_{1})$ can be expressed as the marginal probability of all covering tokens of $x^{n}_{1}$
$P(x^{n}_{1})=\sum_{\vec{t}\in\mathrm{cover}(x^{n}_{1})}P(\vec{t})$ (50)
Proof.
We prove each point as follows:
  1. 1.
    Proof by contradiction, suppose that $\mathcal{S}(\vec{t}_{i})\cap\mathcal{S}(\vec{t}_{j})\neq\emptyset$, then there exists a string $s$ that has two different cover encodings $\vec{t}_{1}$ and $\vec{t}_{2}$. This is impossible since each string $s$ has only one unique encoding.
  2. 2.
    This follows the definition of cover encodings.
Since each $\mathcal{S}(\vec{t})$ is pair-wise disjoint, we arrive at the final equation. We illustrate this Proposition in Figure 7. ∎
Figure 6: Illustration of Proposition H.5 when $|\mathrm{cover}(x^{n}_{1})|{=}3$.
Finally, we prove that BPC extracts all the cover encodings of $x^{n}_{1}$. Proposition H.6 shows the correctness of line 9 in the algorithm, where suppose that the last token starts from $x_{i+1}$, then $\mathrm{encode}(x^{i}_{1})$ must be the encoding before that last token. Since each cover encoding $\vec{t}$ must have a last token cover a suffix within $x^{n}_{1}$, iterating over all positions from $1$ to $n$ guarantees that we extract all possible encodings.
Proposition H.6.
Let $\vec{t}\in\mathrm{cover}(x^{N}_{1})$ and $k=|\vec{t}|$ is the number of tokens in $\vec{t}$. Suppose that $x^{N}_{j+1}$ is a prefix of the $t_{k}$ for some $0\leq j\leq N$, i.e. $x^{N}_{j}\in\mathrm{prefix}(\vec{t}_{k})$, then $t^{k-1}_{1}=\mathrm{encode}(x^{j}_{1})$.
Proof.
Since $\vec{t}=t^{k}_{1}\in\mathrm{cover}(x^{N}_{1})$, then $t^{k-1}_{1}$ must be a valid encoding. As a result, we must have $t^{k-1}_{1}=\mathrm{encode}(x^{j}_{1})$. ∎
Refactoring. Unlike the case for MPE, identifying the case when $P(x^{N}_{n+1}|t^{i}_{1})=P(x^{N}_{n+1}|x^{n}_{1})$ is nontrivial in general for BPE. Nevertheless, the refactoring step aims to reduce the computational complexity, and in general, we can still compute $P(x^{N}_{n+1}|x^{n}_{1})$ by refactoring.
$P(x^{N}_{n+1}|x^{n}_{1})=\frac{P(x^{N}_{1})}{P(x^{n}_{1})},$ (51)
and we use the BPC algorithm to compute $P(x^{N}_{1})$ and $P(x^{n}_{1})$ respectively. Note that this is equivalent to assuming $t_{1}$ is a ${<}\mathrm{start}{>}$ token within $\mathcal{V}^{*}$ (and consider $P(x^{N}_{2}|t_{1}),P(x^{n}_{2}|t_{1})$ instead of $P(x^{N}_{1}),P(x^{n}_{1})$). When pretokenization is used, e.g. split by white spaces, we can identify when $P(x^{N}_{n+1}|t^{i}_{1})=P(x^{N}_{n+1}|x^{n}_{1})$ by using the pretokenization pattern.
H.4 Experiments
[…]